OpenStudy (anonymous):

Anti derivative question:

OpenStudy (anonymous):

I suppose you want to find the antiderivative with respect to x \[I=-\frac{ 1 }{ 12x^4 }\] you can always check it by deriving of course, this is the indefinite integral

OpenStudy (anonymous):



OpenStudy (anonymous):

let me rephrase the question: \[f(x)=\frac{ 1 }{ 3x^3 }\] State is the following is true: 1. \[\int\limits_{0}^{1} f(x) dx=0.5\] 2. the function \[F(x)=3-[\frac{ 1 }{ 6x^2 }]\]

OpenStudy (anonymous):

if G(x) is the anti derivative of f(x) then G'(1)=\[2\frac{ 5}{ 6 }\] in the range (0,infinity)

OpenStudy (anonymous):

3*

OpenStudy (anonymous):

if you want the definite integral: \[\lim_{b \rightarrow \infty}(-\frac{ 1 }{ 12b^4 }+\frac{ 1 }{ 12})=\frac{ 1 }{ 12 }\] P.S: The definite integral should have been: \[-\frac{ 1 }{ 12x^4 }+C\]

OpenStudy (anonymous):

@smarty20two how did you get to -1/12b^4 ? Can you break it up for me in stages? thanks.

OpenStudy (anonymous):

-1/12x^4 could be written as -(1/12)(x^-4) you can keep the constant (-1/12) out and integrate x^-4 you get (-1/12)((x^(-4+1))/(-4+1))+C